# Understanding Why Label Smoothing Degrades Selective Classification and How to Fix It

### v1

@@ 2.4.1 Models and Training
The only parameter varied between training runs is the level of label smoothing $\alpha$. Moreover, in order to isolate the effects of label smoothing, we train all our models from scratch using simple recipes. We purposely avoid popular augmentations such as MixUp [87] and CutMix [86] as these directly affect the training labels, which would interfere with our experiments. For image classification, we train ResNet-50 [27] and ViT-S-16 [13] on ImageNet-1k [63] using only random resized cropping and flipping for data augmentation. To achieve decent accuracy for ViT training from scratch without advanced augmentations, we use sharpness-aware minimisation (SAM) [16, 8]. For semantic segmentation, we train DeepLabV3+ [7] (ResNet-101 backbone) on Cityscapes [12]. We only use random cropping, flipping and colour jitter for augmentations. Full training details for reproducing our results can be found in the Appendix 0.B.
3 Label Smoothing Degrades Selective Classification Performance
@@
Figure 4: Risk-coverage plots for ViT-S-16-SAM trained on ImageNet-1k and DeepLabV3+ (ResNet-101) trained on Cityscapes, for different levels of LS $\alpha$. Although LS may improve error rate at 100% coverage, increasing $\alpha$ consistently degrades selective classification performance.
Figure 5: A few illustrative examples of overconfident errors performed by our LS-trained ResNet-50. Even though $P_{\text{error}}$ is high in all cases the model has predicted low-ranking (amongst all samples) uncertainty. Note that even though the CE model is wrong as well, it has assigned higher ranking uncertainties, reflecting its superior SC ability shown in Fig. 1. As an aside we note the high aleatoric uncertainty in ImageNet, arising from images containing multiple possible labels.
Figure 6: Visualisation of the overconfidence of DeepLabV3+ (ResNet-101) when trained using LS. We see that compared to the model trained using CE, the LS model predicts much lower uncertainty on incorrect ✗ segmentations. In particular, for the erroneous region on the left where the model has predicted parts of the “sidewalk” as “road”, the LS model is highly overconfident. This could have dire consequences in a safety-critical application such as autonomous driving.
Recently, large-scale empirical evaluations of SC for pre-trained models from open-source repositories [20, 4] have revealed that many strong classifiers surprisingly underperform on SC. In this section, we present LS as one possible cause of these previously unexplained results, and explain how LS may affect a model’s ability to perform SC.
To examine the effects of LS on SC performance, we plot RC curves for different architectures and tasks in Figs. 1 and 4. We use $U=-\text{MSP}$ and vary only the LS level $\alpha$ between training runs. We see that training with LS leads to a noticeable degradation for selective classification, despite improvements in top-1 error rate (risk at 100% coverage). That is to say, LS weakens the MSP score’s ability to differentiate  vs . We provide illustrative examples of this overconfidence on ImageNet in Fig. 5. The degradation in performance is especially evident for the low-risk regime, which is relevant to safety-critical scenarios where tolerances for error are low. For example, if the target is to achieve only 1% error on ImageNet, then none of our LS models are able to achieve a meaningful level of coverage, effectively not being able to serve any predictions.
@@
To further highlight the overconfidence caused by LS, we visualise the uncertainty ranking of incorrect pixels for the segmentation of a Cityscapes scene in Fig. 6. We see that the LS-trained model is extremely confident for an erroneous region where it has predicted the “sidewalk” as “road”. Whilst the CE-trained model has made the same error, it is much more uncertain. This illustrates potential danger in an autonomous driving scenario if the vehicle is making decisions based on uncertainty estimates.
@@
Upon inspection, many of the underperforming models benchmarked in [20, 4] are indeed trained using LS, [^3] We include a more detailed discussion of existing benchmarks in Sec. 0.D.2. further supporting our own results. The models are sourced from repositories such as from torchvision [61] and timm [79], where the original training recipes are mainly optimised for top-1 accuracy. [^4] This blog post is a good example https://pytorch.org/blog/how-to-train-state-of-the-art-models-using-torchvision-latest-primitives/ . We note that in the case of ImageNet, LS is such a common technique that it is often used by default, and not even mentioned in papers, e.g. [72, 73]. Overall, our results highlight that blindly optimising for accuracy may result in negative downstream consequences and that practitioners of selective classification need to be aware of the effects of their training recipes.
3.1 Considering Overconfidence Arising from Epistemic Uncertainty
It might intuitively seem odd that LS leads to such degraded SC performance since the transform applied to the true targets in Eq. 9 $(1-\alpha)\bar{\boldsymbol{\pi}}(\boldsymbol{x})+\boldsymbol{u}$ does not change the relative ranking of the maximum probability. However, this is only true for low epistemic uncertainty, i.e., when the model is knowledgeable about the data, so $\boldsymbol{\pi}(\boldsymbol{x};\boldsymbol{\theta})\approx(1-\alpha)\bar{\boldsymbol{\pi}}(\boldsymbol{x})+\boldsymbol{u}$. For such samples, the model has learnt to be more uncertain. However, we cannot assume that epistemic uncertainty is always low. Moreover, we know it leads to both errors and overconfidence (Fig. 3 (right)). One clue that points to epistemic uncertainty is the fact that for the LS-trained models, Figs. 1 and 4 show that many samples have $\pi_{\text{max}}>(1-\alpha)+\alpha/K$, and it is for these such samples that SC performance is degraded the most. As the training targets never exceed $(1-\alpha)+\alpha/K$, this inaccuracy in predicting the smoothed targets could be due to epistemic uncertainty.
@@ 3.2 Comparing Logit Gradients Between CE and LS
To better understand how LS could lead to the poor SC performance exhibited in the RC curves of Figs. 1 and 4, we consider how LS affects logit-level training gradients. These are the first term in the chain rule for backpropagation and so directly contribute to all parameter gradients during training. We take the gradient of $\mathcal{L}^{\text{true}}$ (Eqs. 4 and 9), which the empirical loss approximates, for a single sample,
@@
where in a slight abuse of notation we omit the outer expectation over $p_{\text{data}}(\boldsymbol{x})$ for convenience. We can then define the regularisation gradient on the logits,
@@
which is the difference between the LS and CE gradients. This intuitively represents how LS regularises training at the logit level in comparison to CE. Crucially, it only depends on the true distribution $\bar{\boldsymbol{\pi}}$. Gradient descent involves updating weights in the opposite direction to the gradient. Hence $\alpha\bar{\pi}_{k}$ penalises $v_{k}$ when the true probability $\bar{\pi}_{k}$ is higher. The second term $-\alpha/K$ uniformly increases the logits. This does not affect the softmax output as it is invariant to uniform logit shifts $\boldsymbol{\pi}(\boldsymbol{v})=\boldsymbol{\pi}(\boldsymbol{v}+\eta\boldsymbol{1})$, $\eta\in\mathbb{R}$ [44].
3.3 Label Smoothing Exacerbates Overconfidence and Underconfidence
Figure 7: Illustration of how the regularisation gradient (Eq. 15), which is the difference between LS and CE gradients, affects the logits differently. When $U$ is lower (sharper softmax), the regularisation on the max logit is lower when epistemic uncertainty has caused overconfidence. When $U$ is higher (flatter softmax), the regularisation is higher when epistemic uncertainty has caused underconfidence. Thus, LS exacerbates both of these issues, degrading the rank ordering of $U$ and diminishing the ability of the softmax to differentiate correct  from incorrect  .
We now consider how the regularisation gradient affects the maximum logit,
@@ 3.3 Label Smoothing Exacerbates Overconfidence and Underconfidence
To further build an intuition, we consider overconfidence and underconfidence as a result of epistemic uncertainty. As illustrated in Fig. 7, when $U$ is low, and a model is overconfident due to epistemic uncertainty (i.e. $P_{\text{error}}$ is high), the regularisation gradient is much weaker compared to when $P_{\text{error}}$ is low for the same $U$. On the other hand, when $U$ is high and a model is underconfident due to epistemic uncertainty (i.e. $P_{\text{error}}$ is low), the regularisation gradient is much stronger compared to when $P_{\text{error}}$ is high for the same $U$. Thus, both overconfidence and underconfidence are exacerbated by label smoothing, degrading the model’s ability to perform selective classification.
To test our analysis, we artificially inject epistemic uncertainty into our ImageNet evaluation using ImageNet-Sketch [76], a dataset containing sketches of each ImageNet class (see Fig. 3 for an example). We combine the 50,889 ImageNet-Sketch images with the 50,000 ImageNet evaluation images and plot the new risk vs coverage for all 100,889 samples in Fig. 8. We find that even though the regularisation of LS still improves the error rate, the degradation in SC from LS is exacerbated at lower coverages, where the CE-trained model maintains much lower risk. At 10% coverage of the combined evaluation set, LS leads to accepting many more errors compared to CE, with an increasing proportion originating from ImageNet-Sketch. This demonstrates that LS exacerbates overconfidence arising from epistemic uncertainty, reducing a model’s ability to reject errors.
@10% coverage of ImageNet + Sketch
ImageNet
Sketch
CE
#samples
8994
1094
#errors
93
80
error rate
1.0
7.3
LS $\alpha=0.1$
#samples
8561
1527
#errors
187
211
error rate
2.2
13.8
LS $\alpha=0.2$
#samples
8171
1917
#errors
244
365
error rate
3.0
19.0
LS $\alpha=0.3$
#samples
7915
2173
#errors
297
541
error rate
3.8
24.9
Figure 8: Left: RC curve for ResNet-50 when evaluation is on the combination of ImageNet and ImageNet-sketch (ViT-S-16 can be found in Appendix 0.A). We see that for low-uncertainty predictions, the degradation caused by LS is exacerbated when epistemic uncertainty is artificially introduced. Right: statistics @10% coverage of the combined evaluation set. As the level of LS $\alpha$ increases, the number of errors increases, especially the number of errors from ImageNet-Sketch. This shows that LS exacerbates model overconfidence arising from epistemic uncertainty.
4 Logit Normalisation Improves the SC of LS-Trained Models
Ideally, we would like to find a way to recover from the degradation caused by LS. A recent study by Cattelan and Silva [4] has shown that logit normalisation can improve the SC performance of many pre-trained models. The logits are normalised by their $p$-norm and then the MSP score $\pi_{\text{max}}$ is replaced by the normalised max logit
$U=-v^{\prime}_{\text{max}},\quad\boldsymbol{v}^{\prime}=\frac{\boldsymbol{v}}{\left\lVert\boldsymbol{v}\right\rVert_{p}}=\boldsymbol{v}/\bigl{(}\sum_{k}v_{k}^{p}\bigr{)}^{\frac{1}{p}}~{},$ (17)
where $p$ is found via grid search on a validation set. We investigate the efficacy of this approach, applying it to our LS-trained models. Fig. 9 (left) shows the SC performance with and without logit normalisation, and we indeed find that applying logit normalisation greatly improves SC performance for models trained using LS, allowing for improved error rate (@100% coverage) without compromising on SC. This is further visualised for semantic segmentation in Fig. 10, where logit normalisation successfully mitigates the previously observed overconfidence caused by LS. However, we also find that logit normalisation does not notably improve the ability of CE-trained models. This aligns with the results reported [4] where certain models do not benefit from logit normalisation so the authors suggest “falling back” to the MSP.
Figure 9: Left: RC curves showing the effect of logit normalisation. Logit normalisation improves SC performance on LS-trained models, but has little effect on CE-trained models. DeepLabV3+ can be found in Appendix 0.A. Right: distribution of the max logit $v_{\text{max}}$ given the MSP $\pi_{\text{max}}$ for correct  and incorrect  predictions separately. We see that $v_{\text{max}}$ is higher for  for the LS-trained model, whilst the distributions are roughly the same for CE. This matches the max logit regularisation described in Eq. 16. We calculate the mean±$\pm$±std. in a 0.05-wide sliding window.
Figure 10: Visualisation of the effect of logit normalisation for semantic segmentation (same scene as Fig. 6). Logit normalisation significantly reduces the overconfidence of the LS-trained model, although it has little effect on the uncertainty of the CE-trained model.
@@ 4 Logit Normalisation Improves the SC of LS-Trained Models
Although Cattelan and Silva [4] validate this approach on a large number of pre-trained models, they provide little explanation as to why it is so effective (or indeed why it isn’t effective sometimes). Let us highlight the interaction between logit normalisation and label smoothing. If we examine Eq. 17, we see that it resembles the softmax. However, in this case, the logits are raised to power $p$ rather than being exponentiated. Crucially, whilst the softmax is invariant to uniform shifts in the logits,
@@
this is not the case for $\boldsymbol{v}^{\prime}$. In fact, we have the following inequality for positive [^5] Although this assumption may not necessarily hold, we find that in practice, $\pi_{\text{max}}$ and $v^{\prime}_{\text{max}}$ are dominated by the larger positive logits. We include additional discussion in Sec. 0.D.1. logits $\boldsymbol{v}$:
@@
For all strictly positive vectors $\boldsymbol{v}\in\left(\mathbb{R}_{>0}\right)^{K}$, containing at least two different values, and $p\in[1,+\infty[$, the ratio of the $\infty$-norm and the $p$-norm strictly decreases when summing $v$ with any uniform vector $\eta\boldsymbol{1}$, $\eta$ strictly positive:
$\frac{||\boldsymbol{v}||_{\infty}}{||\boldsymbol{v}||_{p}}>\frac{||\boldsymbol{v}+\eta\boldsymbol{1}||_{\infty}}{||\boldsymbol{v}+\eta\boldsymbol{1}||_{p}}~{}.$ (19)
The proof can be found in Appendix 0.C. Recalling that $||\boldsymbol{v}||_{\infty}=v_{\text{max}}$, Result 1 thus implies that for a given softmax output $\boldsymbol{\pi}(\boldsymbol{v})=\boldsymbol{\pi}(\boldsymbol{v}+\eta\boldsymbol{1})$ and arbitrary corresponding $\boldsymbol{v}$, the greater the value of $\eta$ and thus $v_{\text{max}}$, the lower the value of $v^{\prime}_{\text{max}}$. That is to say logit normalisation increases the estimated uncertainty when the max logit $v_{\text{max}}$ is higher for the same softmax probabilities $\boldsymbol{\pi}$.
Recall that in Sec. 3.3 we found that the maximum logit $v_{\text{max}}$ is less regularised the more likely a prediction is wrong ✗, i.e. when $P_{\text{error}}$ is high (Eq. 16). We plot $v_{\text{max}}$ given the MSP, separately for correct ✓ and incorrect ✗ predictions in Fig. 9, and see an interesting behaviour emerge. For the model trained with LS, the imbalanced logit regularisation described in Eq. 16 has caused $v_{\text{max}}$ to become higher for misclassifications ✗ than correct predictions ✓, i.e. information about $P_{\text{error}}$ has been encoded in $v_{\text{max}}$.
As Result 1 says that for the same softmax output $\boldsymbol{\pi}$, $v^{\prime}_{\text{max}}$ will decrease as $v_{\text{max}}$ increases, logit normalisation will increase uncertainty when $v_{\text{max}}$ is higher, effectively reversing the effect of the imbalanced logit regularisation. This will then improve the rank ordering of the uncertainty score $U$ and thus improve the SC performance of the LS-trained model. Given our analysis and empirical results, we strongly recommend the use of logit normalisation for LS-trained models for performing selective classification.
On the other hand, Fig. 9 (right) also shows that the distributions of $v_{\text{max}}$ given $\pi_{\text{max}}$ for correct ✓ and incorrect ✗ predictions are very similar for the CE-trained model. This explains why, for the models trained using CE, logit-normalisation does not seem to help, as $v_{\text{max}}$ does not provide useful information about $P_{\text{error}}$ given $\pi_{\text{max}}$.

### v2

@@ 3 The effect of label smoothing on selective classification
Experimental details. We investigate large-scale [^2] We additionally provide $32\times 32$ small-scale CIFAR experiments in Sec. A.3. image classification on ImageNet-1k (Russakovsky et al., 2015) and semantic segmentation (pixel-level classification) on Cityscapes (Cordts et al., 2016). For ImageNet, we evaluate on the original validation set and randomly split 50,000 images from the training set for validation. For Cityscapes, we randomly split the original validation set into 100 validation and 400 evaluation images. To estimate the risk, we extract the model output prior to the final interpolation and subsample 5000 labelled pixels per image at random. We evaluate on the same pixels between models. The only parameter varied between training runs is the level of LS $\alpha$, and to isolate the effects of LS, we train all models from scratch using simple recipes. We purposely avoid augmentations such as MixUp (Zhang et al., 2018) and CutMix (Yun et al., 2019) as these directly affect the training labels, which would interfere with our experiments. For ImageNet classification, we train ResNet-50 (He et al., 2016) and ViT-S-16 (Dosovitskiy et al., 2021) using only random resized cropping and flipping for data augmentation. To achieve decent accuracy for ViT training from scratch without advanced augmentations, we use sharpness-aware minimisation (SAM) (Foret et al., 2021; Chen et al., 2022). For semantic segmentation on Cityscapes, we train DeepLabV3+ (Chen et al., 2018) (ResNet-101 backbone) using only random cropping, flipping and colour jitter for augmentations. Full training details for reproducibility can be found in the Appendix B. We will release our code in the future.
3.1 Label smoothing degrades selective classification
@@
Figure 3: Risk-coverage plots for different levels of LS $\alpha$ for different models and tasks (ImageNet classification and Cityscapes semantic segmentation). Although LS may improve error rate at 100% coverage, increasing $\alpha$ increasingly and consistently degrades SC performance.
@@
To examine the effects of LS on SC, we plot RC curves in Fig. 3. We use $U=-\text{MSP}$ and vary only the LS level $\alpha$ between training runs. We see that training with LS leads to a noticeable degradation for selective classification, with higher $\alpha$ worsening the effect. We provide illustrative examples of LS overconfidence on ImageNet in Fig. 9. Although LS leads to slightly better risk at higher coverage, it quickly becomes worse than CE as coverage is reduced. The degradation is especially evident for the low-risk regime, which is relevant to safety-critical scenarios where tolerances for error are strict. For example, if the target is to achieve only 1% error/risk on ImageNet, then all of our LS models have close to zero coverage, rendering them effectively useless (Figs. 1 and 3). To further highlight the overconfidence caused by LS, we visualise the uncertainty ranking of incorrect pixels for the segmentation of a Cityscapes scene in Fig. 1. The LS-trained model is extremely confident for an erroneous region where it has predicted the “sidewalk” as “road”. Whilst the CE model has made the same error, it is much more uncertain. This illustrates potential danger in an autonomous driving scenario if the vehicle is making decisions based on uncertainty estimates.
@@
Upon inspection, many of the SC-underperforming models benchmarked in (Galil et al., 2023; Cattelan & Silva, 2024) are trained using LS, aligning with our results. The models are sourced from repositories such as torchvision (Paszke et al., 2019) and timm (Wightman, 2019), where the training recipes are optimised for top-1 accuracy. We note that for ImageNet, LS is such a common technique that it is often used by default, and not even mentioned in papers, e.g. (Tan & Le, 2019; 2021). Overall, our results highlight that only optimising for accuracy may result in negative downstream consequences and that practitioners of SC need to be aware of the effects of their training recipes. We include additional discussion of existing benchmarks and model checkpoints in Sec. E.4.
4 Towards explaining the negative effect of LS on SC
In the following section, we attempt to delve deeper into why our empirical results (and those in (Zhu et al., 2024)) occur, shedding light on a previously unexplained phenomenon. It might intuitively seem odd that LS degrades SC performance as the transform applied to the true targets in Eq. 8 $(1-\alpha)\bar{\boldsymbol{\pi}}(\boldsymbol{x})+\alpha\boldsymbol{u}$ reduces the max probability for all samples but does not change the relative ranking. However, this only carries over to the model when it is knowledgeable about the data, so it is well fit $\boldsymbol{\pi}(\boldsymbol{x};\boldsymbol{\theta})\approx(1-\alpha)\bar{\boldsymbol{\pi}}(\boldsymbol{x})+\alpha\boldsymbol{u}$. For such samples, the model has learnt to be more uncertain. However, we cannot assume that the model is equally well fit on all regions of the data distribution. [^3] For the sake of simplicity we avoid the term epistemic uncertainty (see Sec. E.2 for brief discussion).
Through analysing the logit gradients, we provide a potential explanation: LS suppresses the max logit differently depending on how well fit the model is to a given sample – suppression is stronger(weaker) the more likely a model is correct(wrong), harming the model’s ability to separate ✓ vs ✗.
@@ 4.1 Comparing logit gradients between CE and LS
To better understand how LS could lead to degraded SC, we consider how LS affects logit-level training gradients. These are the first term in the chain rule for backpropagation and so directly contribute to all parameter gradients during training. We take the gradient of $\mathcal{L}^{\text{true}}$ (Eqs. 4 and 8), [^4] Here we assume that the empirical loss $\mathcal{L}$ (Eqs. 2 and 6) approximates $\mathcal{L}^{\text{true}}$, in order to relate our discussion to $P_{\text{error}}$. However, the same analysis on the empirical loss leads to similar conclusions (see Appendix C).
@@
for a single sample, where in a slight abuse of notation we omit the outer expectation over $p_{\text{data}}(\boldsymbol{x})$ for convenience. We can then define the regularisation gradient on the logits,
@@
which is the difference between the LS and CE gradients. This represents how LS influences training at the logit level in comparison to CE. Notably, it only depends on the target $\bar{\boldsymbol{\pi}}$. Gradient descent involves updating weights in the opposite direction to the gradient. Hence $\alpha\bar{\pi}_{k}$ suppresses $v_{k}$ when the true probability $\bar{\pi}_{k}$ is higher. The second term $-\alpha/K$ uniformly increases the logits for all samples, which does not affect the softmax as it is invariant to uniform logit shifts $\boldsymbol{\pi}(\boldsymbol{v})=\boldsymbol{\pi}(\boldsymbol{v}+\eta\boldsymbol{1})$, $\eta\in\mathbb{R}$.
4.2 Imbalanced regularisation degrades uncertainty ranking of  vs
Figure 4: How the regularisation gradient (Eq. 14), which is the difference between LS and CE gradients, affects the logits differently. LS affects the max logit differently depending on how well fit the model is for a given sample $\boldsymbol{x}$. On the left when $U$ is lower (sharper softmax), the regularisation on the max logit is lower when the model is poorly fit and likely to be wrong. On the right when $U$ is higher (flatter softmax), the regularisation is higher when the model is poorly fit and more likely to be correct. Thus, LS degrades the softmax’s ability to separate  vs  , hurting SC.
We now consider how the regularisation gradient affects the maximum logit $v_{\text{max}}$,
@@ 4.2 Imbalanced regularisation degrades uncertainty ranking of  vs
which shows that the regularisation on $v_{\text{max}}$ decreases as the probability of error increases. This directly impacts softmax-based $U$ such as MSP (see Eq. 1), especially for lower uncertainties where $v_{\text{max}}$ dominates the softmax. Eq. 15 suggests that for predictions that share the same value of $U$, those with higher $P_{\text{error}}$ (more likely ✗), will have $v_{\text{max}}$ regularised less, whilst predictions with lower $P_{\text{error}}$ (more likely ✓) will have $v_{\text{max}}$ regularised more. Thus, softmax-based $U$ will have the relative ranking of correct ✓ and incorrect ✗ predictions degraded – less uncertain on ✗ , more uncertain on ✓. To further build an intuition, we consider how a model may not be uniformly well-fit over the data distribution during training. As illustrated in Fig. 4, varying levels of fit with regards to the true distribution $\bar{\pi}$ leads to differing $P_{\text{error}}$ for predictions with similar $U$, resulting in the degradation described above. Finally, as the regularisation gradient in Eq. 15 is proportional to $\alpha$, this explains why stronger LS leads to greater degradation. In summary: for a given predicted $U$ the $P_{\text{error}}$ will vary depending on how well fit the model is. LS suppresses the max logit more(less) when $P_{\text{error}}$ is lower(higher), degrading the ranking of  vs  for softmax-based $U$, hurting SC. We note that although our analysis is performed on training gradients, all our experiments are performed on evaluation data, suggesting that effects indeed generalise to unseen data. We also note that our explanation is purely based on the loss, and thus generalises across network architectures and tasks, matching the empirical results in Sec. 3.
LS empirically leads to higher $v_{\text{max}}$ on misclassifications . To further validate the effects of Eq. 15, we plot the mean±$\pm$±std. of $v_{\text{max}}$ given $\pi_{\text{max}}$ for ResNet-50 on evaluation data in Fig. 5. We see that for LS the distribution of correct ✓  is below incorrect ✗, whilst for CE they are roughly similar. This provides further empirical support for Eq. 15 which states that when using LS, $v_{\text{max}}$ is more strongly suppressed for lower $P_{\text{error}}$. For results on other models see Sec. A.2.
LS empirically leads to increasing overconfidence on less-well-fit data. To further investigate how LS affects uncertainty differently depending on how well-fit/knowledgeable the model is about data, we artificially introduce distribution-shifted data that we expect our models to be worse fit on. We use ImageNet-Sketch (Wang et al., 2019), a dataset containing sketches of each ImageNet class and evaluate on the combination of the 50,889 ImageNet-Sketch and 50,000 ImageNet evaluation images in Fig. 6. Even though the regularisation of LS improves the error rate, the degradation of SC at lower coverages from LS is exacerbated by the distribution-shifted data. At 10% coverage of the combined evaluation set, LS leads to accepting many more errors compared to CE, with an increasing proportion originating from ImageNet-Sketch. That is to say, LS leads to increasing overconfidence on less-well-fit data (ImageNet-Sketch). This empirically highlights the situation illustrated on the left of Fig. 4 where confident and well-fit samples have their max logit suppressed, whilst confident but poorly fit samples do not, leading to samples with higher $P_{\text{error}}$ being relatively less uncertain (overconfident). For results on other models see Sec. A.2.
Figure 5: Distribution of the max logit $v_{\text{max}}$ given the MSP $\pi_{\text{max}}$ for correct  and incorrect  predictions on evaluation data. $v_{\text{max}}$ is lower for  for the LS model, whilst the distributions are roughly similar for CE. This empirically matches the imbalanced max logit regularisation described in Eq. 15. We calculate the mean±$\pm$±std. in a 0.05-wide sliding window.
@10% coverage of ImageNet + Sketch
ImageNet Sketch
CE #samples 8994 1094
#errors 93 80
error rate 1.0 7.3
LS $\alpha=0.1$samples 8561 1527
#errors 187 211
error rate 2.2 13.8
LS $\alpha=0.2$samples 8171 1917
#errors 244 365
error rate 3.0 19.0
LS $\alpha=0.3$samples 7915 2173
#errors 297 541
error rate 3.8 24.9
Figure 6: Left: Evaluation on the combination of ImageNet + ImageNet-sketch. We see that for low-uncertainty predictions, the degradation caused by LS is exacerbated when distribution shift is artificially introduced (vs Fig. 3). Right: statistics @10% coverage of the combined evaluation set. As the level of LS $\alpha$ increases, the number of accepted errors increases, especially from ImageNet-Sketch. This shows that LS leads to increasing overconfidence on less-well-fit data.
5 Logit normalisation improves the SC of LS-trained models
Ideally, we would like to find a way to recover from the degradation caused by LS. A recent empirical study by Cattelan & Silva (2024) has shown that logit normalisation can improve the SC performance of many (but not all) pre-trained models. During deployment the logits are normalised by their $p$-norm and then the MSP score $\pi_{\text{max}}$ is replaced by the normalised max logit,
$U=-v^{\prime}_{\text{max}}=-\max\nolimits_{k}v^{\prime}_{k},\quad\boldsymbol{v}^{\prime}=\boldsymbol{v}/\left\lVert\boldsymbol{v}\right\rVert_{p}=\boldsymbol{v}\big{/}\bigl{(}\sum\nolimits_{k}|v_{k}|^{p}\bigr{)}^{\frac{1}{p}}~{},$ (16)
@@ 5 Logit normalisation improves the SC of LS-trained models
Figure 7: RC curves with inference-time logit normalisation. Logit normalisation improves SC performance on LS models, but has little effect on CE models. For other models, see Sec. A.2.
Figure 8: Visualisation of the effect of logit normalisation for segmentation (same scene as Fig. 1). Logit normalisation significantly reduces the overconfidence of the LS-trained model, although it has little effect on the uncertainty of the CE-trained model. We provide more figures in Appendix F.
@@
Although Cattelan & Silva (2024) empirically validate this approach on a large number of pre-trained models, it remains unclear as to why it is so effective, or why it isn’t effective sometimes (further discussion in Sec. E.5). Let us shed light on the interaction between logit normalisation and LS. If we examine Eq. 16, we see that it resembles the softmax, however, the logits are raised to power $p$ rather than exponentiated. Crucially, whilst the softmax is invariant to uniform shifts in the logits,
@@
this is not the case for $\boldsymbol{v}^{\prime}$. In fact, we have the following inequality for positive [^6] Although this assumption may not necessarily hold, we find that in practice, $\pi_{\text{max}}$ and $v^{\prime}_{\text{max}}$ are dominated by the larger positive logits, meaning the behaviour discussed still occurs empirically (Secs. E.3 and 20). logits $\boldsymbol{v}$:
@@
For all vectors $\boldsymbol{v}\in\left(\mathbb{R}_{>0}\right)^{K}$ with at least two different values and $p\in[1,+\infty[$, the ratio of the $\infty$-norm and $p$-norm strictly decreases when summing $\boldsymbol{v}$ with any uniform vector $\eta\boldsymbol{1}$, $\eta>0$:
$v^{\prime}_{\text{max}}(\boldsymbol{v})=||\boldsymbol{v}||_{\infty}\big{/}||\boldsymbol{v}||_{p}\quad>\quad||\boldsymbol{v}+\eta\boldsymbol{1}||_{\infty}\big{/}||\boldsymbol{v}+\eta\boldsymbol{1}||_{p}=v^{\prime}_{\text{max}}(\boldsymbol{v}+\eta\boldsymbol{1})~{}.$ (18)
The proof can be found in Appendix D. Recalling that $||\boldsymbol{v}||_{\infty}=v_{\text{max}}$, Result 1 thus implies that for a given softmax output $\boldsymbol{\pi}(\boldsymbol{v})=\boldsymbol{\pi}(\boldsymbol{v}+\eta\boldsymbol{1})$ and arbitrary corresponding $\boldsymbol{v}$, the greater the value of $\eta$ and thus $v_{\text{max}}$, the lower the value of the normalised max logit $v^{\prime}_{\text{max}}$. That is to say logit normalisation increases uncertainty when the max logit $v_{\text{max}}$ is higher for the same softmax probabilities $\boldsymbol{\pi}$. Recall how in Sec. 4.2 we show how imbalanced logit regularisation leads to degradation in SC. In particular Fig. 5 shows how LS leads to higher $v_{\text{max}}$ given $\pi_{\text{max}}$ for errors ✗ – this implies that, independent of the value of $\pi_{\text{max}}$, information about $P_{\text{error}}$ has been encoded in $v_{\text{max}}$. We see now that logit normalisation will increase the uncertainty (ranking) of errors ✗  relative to correct predictions ✓  using the information in $v_{\text{max}}$, leading to improved SC for an LS model. That is to say logit normalisation effectively reverses the effect of the imbalanced logit regularisation from LS, improving SC. We note that Cattelan & Silva (2024) find that models that are less confident (on both ✓ and ✗) tend to benefit more from logit normalisation, again pointing towards LS. Given our analysis and empirical results, we strongly recommend logit normalisation for LS-trained models when performing SC. On the other hand, Fig. 5 also shows that the distributions of $v_{\text{max}}$ given $\pi_{\text{max}}$ for correct ✓ and incorrect ✗ predictions are very similar for the CE model. This explains why for CE, logit-normalisation does not seem to help, as $v_{\text{max}}$ does not provide useful information about $P_{\text{error}}$ given $\pi_{\text{max}}$.